# Constraining first-order phase transition inside neutron stars with application of Bayesian techniques on PSR J0437-4715 NICER data

### v2

@@ 1 Introduction
Neutron stars are among the most extreme laboratories for cold, dense matter in the Universe, reaching core densities that can exceed those of atomic nuclei by several factors. Under these extreme conditions, unusual phases of matter—such as deconfined quarks or hyperons—may emerge, possibly leading to first-order phase transitions in the stellar interior. Identifying these transitions and understanding how they modify key observables like neutron star masses, radii, and spin evolution is a central challenge in astrophysics. Precise measurements of mass and radius, therefore, are vital for probing the neutron star equation of state (EOS) and discerning whether such exotic forms of matter exist.
@@
On the theoretical side, Bayesian inference approaches—supplemented by multi-messenger data—have been successful in constraining the likely densities and depth ranges where phase transitions might occur (see eg. Brandes et al. (2023a); Essick et al. (2023); Gorda et al. (2023); Somasundaram et al. (2023); Alford et al. (2013); Komoltsev (2024)). Nonetheless, definitive observational evidence for first-order phase transitions remains elusive. Motivated by this, we apply Bayesian inference to the new NICER data on PSR J0437--4715 to further assess the plausibility of such transitions. In particular, we employ a polytrope-based meta-model parameterization (Hebeler et al., 2013), allowing a systematic exploration of the physical conditions that could trigger a first-order phase transition within the neutron star core.
@@ 2.1 Prior Settings
To describe the equation of state (EOS), we adopt a three-segment polytropic model incorporating a potential first-order phase transition. This transition is characterized by two adjustable parameters: the transition depth ($\Delta\varepsilon$) and the critical transition density ($\varepsilon_{t}$). These parameters enable us to represent phase changes between the first and second polytropic sections and between the second and third sections. Our approach aligns with the methodologies employed by the NICER collaboration and extends the standard polytropic (PP) model, which accommodates phase transitions when the adiabatic index $\gamma$ approaches zero (Hebeler et al., 2013; Rutherford et al., 2024).
@@ 2.2 Implemented Data
×$\times\prod_{j}p\left(M_{j},R_{j}\mid d_{\mathrm{NICER},\mathrm{j}}\right).$
Here, $\theta$ represents the set of parameters describing the neutron star equation of state (EOS), which includes adiabatic indices $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$, as well the transition density $\epsilon_{t}$ and transition depth $\Delta\epsilon$. The symbol $d$ denotes the data set, which consists of the mass and radius measurements of neutron stars, that are inferred from NICER observations. The term $M$ refers to the model used to describe the neutron star EOS, which incorporates both polytropic segments and the possibility of phase transitions. The likelihood function $p(\theta,\epsilon|d,M)$ expresses the probability of observing the data $d$ given the parameters $\theta$ and model $M$, with $\epsilon$ representing the energy density or other relevant quantities depending on the model. The prior distribution $p(\theta|M)$ reflects the initial beliefs about the parameters before considering the data. The term $p(\epsilon|\theta,M)$ is the likelihood of the energy density $\epsilon$ given the parameters $\theta$ and model $M$. Finally, the product $\prod_{j}p(M_{j},R_{j}|d_{\text{NICER},j})$ represents the likelihood of each NICER measurement for different pulsars, where $M_{j}$ and $R_{j}$ are the mass and radius of the $j$-th neutron star, and $d_{\text{NICER},j}$ is the corresponding observed data. We will conduct two different inferences: the first using only PSR J0030+0451 and PSR J0740+6620, and the second including the new observation of PSR J0437--4715. These results will be compared to inferences made with a model that does not include phase transitions. For the without phase transition (NPT) model, each dataset (with and without PSR J0437--4715) will also be analyzed. Comparisons between the same dataset under different models will be made using Bayes factors to determine whether the new observation provides any preference for phase transitions.
@@ 3.1 Constraint on Equation of State parameters
The only difference between these two models is the inclusion of a phase transition. Thus, comparing the Bayesian evidence of these models is meaningful, as it provides crucial information on whether the observations favor the existence of a phase transition. The Bayes factor is the standard method in Bayesian statistics for comparing two models. In this analysis, the Bayes factor for comparing the NPT model with the PT model is defined as $B(\mathrm{NPT}/\mathrm{PT})=\exp\left(\ln Z_{1}-\ln Z_{2}\right)=Z_{1}/Z_{2}$ where $Z_{1}=p(d\mid\mathrm{NPT})$ represents the Bayesian evidence for the NPT model, and $Z_{2}=p(d\mid\mathrm{PT})$ represents the Bayesian evidence for the PT model. In this context,$d$ denotes the observed dataset. Table 2 lists the logarithms of the Bayesian evidence $\ln(Z)$for these two models, denoted as PT and NPT models, under two different observation data sets, without PSR J0437--4715 and this source.
In our analysis, the Bayes’ factor for the scenario that includes PSR J0437--4715 is 2.34, slightly smaller than 3.2, A Bayes’ factor lower than 3.2 is a indicator that there is no any clear evidence favor one model over another. which suggests that current observations, even with the inclusion of new sources, still need to be more precise to distinguish between models with and without phase transitions definitively. However, we are very close to this threshold, and comparing this value with the Bayes’ factor 1.27 without including PSR J0437--471, the improvement from this new observation PSR J0437--4715 appears to be crucial since it could increase the Bayes’ factor significantly. By using Bayes Factor we see that our model shows no clear preference for a polytrope model with or without a phase transition. However, the incorporation of new observational data can increase the Bayes factor in favor of the NPT model. this indicates that PSR J0437--4715 has the potential to identify phase transitions inside neutron stars, especially with its comparably high-precision radius measurement.

### v3

@@ 1 Introduction
Neutron stars are among the most extreme laboratories for cold, dense matter in the Universe, reaching core densities that can exceed those of atomic nuclei by several factors. Under these extreme conditions, unusual phases of matter—such as hyperon or deconfined quarks—may emerge, possibly leading to first-order phase transitions in the stellar interior (Christian et al., 2024). Identifying these transitions and understanding how they modify key observables like neutron star masses, radii, and spin evolution is a central challenge in astrophysics. Precise measurements of mass and radius, therefore, are vital for probing the neutron star equation of state (EOS) and discerning whether such exotic forms of matter exist.
@@
On the theoretical side, Bayesian inference approaches—supplemented by multi-messenger data—have been successful in constraining the likely densities and depth ranges where phase transitions might occur (see eg. Tang et al. (2021); Brandes et al. (2023a); Essick et al. (2023); Gorda et al. (2023); Somasundaram et al. (2023); Alford et al. (2013); Komoltsev (2024)). Nonetheless, definitive observational evidence for first-order phase transitions remains elusive. Motivated by this, we apply Bayesian inference to the new NICER data on PSR J0437--4715 to further assess the plausibility of such transitions. In particular, we employ a polytrope-based meta-model parameterization (Hebeler et al., 2013), allowing a systematic exploration of the physical conditions that could trigger a first-order phase transition within the neutron star core.
@@ 2.1 Prior Settings
To describe the equation of state (EOS), we adopt a three-segment polytropic model incorporating a potential first-order phase transition. This transition is characterized by two adjustable parameters: the transition depth ($\Delta\varepsilon$) and the critical transition density ($\varepsilon_{t}$). These parameters enable us to represent phase changes between the first and second polytropic sections and between the second and third sections. Our approach aligns with the methodologies employed by the NICER collaboration and extends the standard polytropic (PP) model.(Hebeler et al., 2013; Rutherford et al., 2024). Normally the standard PP model contains three segment of polytropes. The fact that the three-segment model is primarily used here does not imply that it is sufficient for investigating questions such as quantitative likelihood differences, see more discussions in (Brandes & Weise, 2025). Since the primary goal of our study is to focus on directly comparing with widely-adopted PP model in the literature and investigate the role of phase transition we did not investigate the PP model with more than three segments.
@@ 2.2 Implemented Data
×$\times\prod_{j}p\left(\mathcal{M}_{j},R_{j}\mid d_{\mathrm{NICER},\mathrm{j}}\right).$
Here, $\bm{\theta}$ represents the set of parameters describing the neutron star equation of state (EOS), which includes adiabatic indices $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$, as well the transition density $\varepsilon_{t}$ and transition depth $\Delta\varepsilon$. The symbol $\bm{d}$ denotes the data set, which consists of the mass and radius measurements of neutron stars, that are inferred from NICER observations. The term $\mathcal{M}$ refers to the model used to describe the neutron star EOS, which incorporates both polytropic segments and the possibility of phase transitions. The likelihood function $p(\bm{\theta},\varepsilon|\bm{d},\mathcal{M})$ expresses the probability of observing the data $\bm{d}$ given the parameters $\bm{\theta}$ and model $\mathcal{M}$, with $\varepsilon$ representing the energy density or other relevant quantities depending on the model. The prior distribution $p(\bm{\theta}|\mathcal{M})$ reflects the initial beliefs about the parameters before considering the data. The term $p(\varepsilon|\bm{\theta},\mathcal{M})$ is the likelihood of the energy density $\varepsilon$ given the parameters $\bm{\theta}$ and model $\mathcal{M}$. Finally, the product $\prod_{j}p(\mathcal{M}_{j},R_{j}|d_{\text{NICER},j})$ represents the likelihood of each NICER measurement for different pulsars, where $\mathcal{M}_{j}$ and $R_{j}$ are the mass and radius of the $j$-th neutron star, and $d_{\text{NICER},j}$ is the corresponding observed data. We will conduct two different inferences: the first using only PSR J0030+0451 and PSR J0740+6620, and the second including the new observation of PSR J0437--4715. These results will be compared to inferences made with a model that does not include phase transitions. For the without phase transition (NPT) model, each dataset (with and without PSR J0437--4715) will also be analyzed. Comparisons between the same dataset under different models will be made using Bayes factors to determine whether the new observation provides any preference for phase transitions.
@@ 3.1 Constraint on Equation of State parameters
The only difference between these two models is the inclusion of a phase transition. Thus, comparing the Bayesian evidence of these models is meaningful, as it provides crucial information on whether the observations favor the existence of a phase transition. The Bayes factor is the standard method in Bayesian statistics for comparing two models. In this analysis, the Bayes factor for comparing the NPT model with the PT model is defined as $B(\mathrm{NPT}/\mathrm{PT})=\exp\left(\ln Z_{1}-\ln Z_{2}\right)=Z_{1}/Z_{2}$ where $Z_{1}=p(\bm{d}\mid\mathrm{NPT})$ represents the Bayesian evidence for the NPT model, and $Z_{2}=p(\bm{d}\mid\mathrm{PT})$ represents the Bayesian evidence for the PT model. In this context,$\bm{d}$ denotes the observed dataset. Table 2 lists the logarithms of the Bayesian evidence $\ln(Z)$for these two models, denoted as PT and NPT models, under two different observation data sets, without PSR J0437--4715 and this source.
In our analysis, the Bayes’ factor for the scenario that includes PSR J0437--4715 is 2.34, slightly smaller than 3.2, A Bayes’ factor lower than 3.2 indicates that there is no clear evidence favoring one model over another. This suggests that current observations, even with the inclusion of new sources, still need to be more precise to distinguish between models with and without phase transitions definitively. However, we are very close to this threshold, and comparing this value with the Bayes’ factor 1.40 without including PSR J0437--471, the improvement from this new observation PSR J0437--4715 appears to be crucial since it could increase the Bayes’ factor significantly. By using Bayes Factor we see that our model shows that there is no clear preference for a polytrope model with or without a phase transition. However, the incorporation of new observational data can increase the Bayes factor in favor of the NPT model. this indicates that PSR J0437--4715 has the potential to identify phase transitions inside neutron stars, especially with its comparably high-precision radius measurement.